One pound of air R = (53.3 ft-lbf/lbm-°R) is compressed isothermally in a closed system from 15 psia and 80°F to 90 psia. Compute the heat transfer.


1
Expert's answer
2021-02-22T07:31:33-0500

As we know that isothermal process is constant temperature process. As we know that internal energy is function of temperature so change in internal energy will be zero.

As per first law of thermodynamics

"\Delta Q= \Delta U+ \Delta W"

"\Delta U= 0" ,

and we know that, work done in isothermal process

"\Delta W= 2.303RT(log(\frac{P_1}{P_2})"

and as per first law of thermodynamics

"\Delta Q= 2.303RT(log(\frac{P_1}{P_2})"

"\Delta Q= 2.303\times 53.3\times80 (log(\frac{15}{90})"

"\Delta Q=-7641.44" ft-lb"_f" , is the heat transfer